### v2

# Projected Subgradient Ascent for Convex Maximization

Abstract.
We consider the problem of maximizing a convex function over a closed convex set in a real Hilbert space. For linear functions, we show that a single orthogonal projection suffices to obtain an approximate solution. For continuous convex functions over convex sets, we show that projected subgradient ascent converges to a first-order stationary point when using arbitrarily large step sizes. Taking the step sizes to infinity leads to a deterministic variant of the conditional gradient algorithm, and iterated linear optimization as a special case.
Keywords: Linear optimization, Orthogonal projection, Nonlinear programming, Subgradient methods.
MSC: 90C05, 90C22, 90C25
1. Introduction
We consider the problem of maximizing a convex function over a nonempty closed and convex set $S$ in a real Hilbert space $\mathcal{H}$. Of particular interest is the case of linear optimization:
(1) $\max_{x\in S}\;\langle c,x\rangle$
This problem can be solved by various methods, including but not limited to interior-point methods, the simplex algorithm, proximal methods, and projected gradient ascent (e.g., [18, 1, 6, 8, 11]).
Here we show that one orthogonal projection suffices to obtain an approximate solution to a linear optimization problem. More generally, the approach can be viewed as a single step of projected gradient ascent using a large step size.
Let $P_{S}(x)$ denote the unique point in $S$ closest to $x$, i.e., the orthogonal projection. We show that $P_{S}(x_{0}+\eta c)$ converges to the unique optimal solution of (1) closest to $x_{0}$ as $\eta\to\infty$ (Theorem 2.7). We also give tight bounds on the objective value of the solution obtained using finite $\eta$ (Lemma 2.4).
Beyond linear objectives, we investigate the use of projected subgradient ascent for maximizing convex functions. In the convex minimization setting, projected subgradient descent requires vanishing step sizes. In contrast, we show that for convex maximization, projected subgradient ascent converges to a first-order stationary point when using arbitrarily large step sizes (Theorem 3.14). This result holds for continuous convex functions in infinite-dimensional Hilbert spaces. Previous work on nonconvex minimization in finite-dimensional spaces (e.g., [2]) typically rely on differentiability, Lipschitz continuity of the gradient, and the Kurdyka-Łojasiewicz inequality to ensure convergence. Our analysis removes these structural requirements by leveraging convexity of the function being maximized.
In practice, the large step size regime of projected subgradient ascent may lead to faster convergence and yield meaningful behavior. In the limit as the step size goes to infinity, projected subgradient ascent leads to the conditional gradient method with unit step size (Section 3.3). This limit also generalizes the iterated linear optimization paradigm introduced in [13].
The idea of using a single projection for linear optimization was previously considered in [15] in the context of linear programming where $S$ is a polytope. The same procedure was rediscovered in [19] under the assumption of strict complementarity, and subsequently extended to infinite-dimensional convex problems satisfying a sharpness property in [9]. Here we consider general convex sets that may have smooth boundaries. Such sets arise in a variety of settings including in semidefinite programming. The idea of using a single projection for linear optimization was also considered recently, and independently of our work, in [22]. Compared to [22], the bounds we derive in Section 2 are tighter, and we prove convergence to a unique optimum solution besides proving convergence of the objective value as was done in [22].
The reduction of linear optimization to orthogonal projection (Section 2) can be used both to understand the relative complexity of the two operations and to derive new algorithms for linear optimization using existing algorithms for projection.
[10] undertook a complexity analysis and show that on many domains—the simplex, $\ell_{p}$-balls, nuclear-norm ball, the flow polytope, Birkhoff polytope, and permutahedron—the best known algorithms for linear optimization are asymptotically faster than the best known algorithms for orthogonal projection. The reduction of linear optimization to orthogonal projection gives further evidence that linear optimization over a convex set is no harder than projecting to the same set. While [10] and [22] suggest this as a negative result for projection-based methods, we emphasize that the reduction can be used to obtain positive results, in that efficient algorithms for projections can lead to efficient algorithms for linear optimization.
2. Single Projection for Linear Optimization
Let $\mathcal{H}$ be a real Hilbert space and $S\subseteq\mathcal{H}$ be a nonempty closed and convex subset. Throughout the paper, we assume that the maximum in (1) is attained by at least one point of $S$. If $S$ is weakly compact, then the maximum is always attained. Define the set of maximizers
$\mathcal{M}(c)=\operatorname*{\arg\!\max}_{x\in S}\,\langle c,x\rangle.$
Since $S$ is closed and convex, and $\langle c,x\rangle$ is a linear function of $x$, the set of maximizers $\mathcal{M}(c)$ is non-empty, closed, and convex. Let $x_{0}\in\mathcal{H}$. Denote by $\|\cdot\|$ the norm induced by the inner product. By the Hilbert Projection Theorem, there is a unique element $x^{*}\in\mathcal{M}(c)$ closest to $x_{0}$,
$x^{*}=\operatorname*{\arg\!\min}_{x\in\mathcal{M}(c)}\|x-x_{0}\|.$
Consider the orthogonal projection map $P_{S}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to S$ taking a point $x\in\mathcal{H}$ to the unique closest point of $x$ in $S$,
$P_{S}(x)=\operatorname*{\arg\!\min}_{y\in S}\|y-x\|.$
For $\eta\in\mathbb{R}$, let
$x^{\eta}=P_{S}(x_{0}+\eta c).$
[…]
Figure 1 illustrates the convergence in the case of an ellipse in the plane with $x_{0}$ at the origin. Since $x^{\eta}$ is the orthogonal projection of $\eta c$ we have that $\eta c-x^{\eta}$ is perpendicular to the ellipse at $x^{\eta}$. Therefore $x^{\eta}$ maximizes $\langle\eta c-x^{\eta},x\rangle$. As $\eta$ grows, $\eta c-x^{\eta}$ becomes parallel to $c$, and in the limit $x^{\eta}$ maximizes $\langle c,x\rangle$.
Figure 1. Linear maximization of $\langle c,x\rangle$ via a single projection $x^{\eta}=P_{S}(c\eta)$ with $\eta\to\infty$.
The use of a single projection for linear optimization was considered in [15] and [19] in the context of linear programming, where projecting a suitably rescaled cost vector yields an optimal basic feasible solution. [9] extended these results to infinite-dimensional spaces under a sharpness assumption. While these work sought an explicit $\eta$ that produces an exact solution in the polyhedral setting, we consider more general (non-polyhedral) closed convex sets in a Hilbert space and derive bounds on $\eta$ that ensure an arbitrarily close approximation to the objective value. Note that for smooth sets, no finite $\eta$ achieves optimality.
A result similar to Lemma 2.4 appears in [22]. However, our lemma provides a more precise characterization that depends on $x_{0}$. This also allows us to prove convergence of $P_{S}(x_{0}+\eta c)$ in Theorem 2.7, including in the case of general (non-polyhedral) convex sets $S$.
Practical use of $x^{\eta}$ to approximate $x^{*}$ requires a choice for $\eta$. The choice should balance the computational complexity of computing the projection $x^{\eta}=P_{S}(x_{0}+\eta c)$ and the quality of the approximation. For some convex sets, projection can be performed efficiently. Methods based on alternating projections, such as Dykstra’s algorithm [12], can also be used in various settings.
The quality of an approximate solution can be measured in terms of the difference in objective value $\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle$. We first provide a bound on the value of $\eta$ sufficient for some desired approximation. Then we demonstrate convergence of the solution.
We start by recalling some basic results.
Definition 2.1.
(Normal Cone) For a point $x\in S$, the normal cone of $S$ at $x$ is the set
$N_{S}(x)=\{y\in\mathcal{H}\mathrel{\mathop{\ordinarycolon}}\langle y,x\rangle\geq\langle y,z\rangle\,\forall z\in S\}.$
The following two lemmas follow from the definition of normal cones and the optimality condition for projections, i.e., $\nabla f(y)^{T}(z-y)\geq 0$ for all $z\in S$ where $y=P_{S}(x)$ and $f(y)=\tfrac{1}{2}\|y-x\|^{2}$.
Lemma 2.2.
Let $x\in\mathcal{H}$ and $y\in S$. Then
$y\in\mathcal{M}(x)\text{ if and only if }x\in N_{S}(y).$
Lemma 2.3.
Let $x\in\mathcal{H}$ and $y\in S$. Then
$y=P_{S}(x)\text{ if and only if }x-y\in N_{S}(y).$
The next key lemma shows $\langle c,x^{\eta}\rangle$ approaches $\langle c,x^{*}\rangle$ as $\eta$ grows.
Lemma 2.4.
Let $\eta>0$. Then
$0\leq\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle\leq\frac{\|x^{*}-x_{0}\|^{2}-\|x^{\eta}-x_{0}\|^{2}}{2\eta}.$
Proof.
Since $x^{\eta}$ is the closest point in $S$ from $x_{0}+\eta c$,
$\|(x_{0}+\eta c)-x^{\eta}\|^{2}\leq\|(x_{0}+\eta c)-x^{*}\|^{2}.$
Expanding and rearranging the terms,
$2\eta\left(\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle\right)\leq\|x^{*}-x_{0}\|^{2}-\|x^{\eta}-x_{0}\|^{2}.$
The left-hand side is nonnegative because $x^{*}$ maximizes $\langle c,x\rangle$. We obtain the desired inequality by dividing both sides by $2\eta$. ∎
When $S$ is bounded and $x_{0}\in S$, we can bound the approximation error using the diameter of $S$.
Observation 2.5.
Let $x_{0}\in S$ and $\eta>0$. Then
$0\leq\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle\leq\frac{\mathrm{diam}(S)^{2}}{2\eta}.$
Observation 2.5 implies we can choose $\eta=\mathrm{diam}(S)^{2}/(2\epsilon)$ to ensure $\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle\leq\epsilon$. The following example illustrates that this does not mean that $x^{\eta}$ is close to $x^{*}$. In fact $x^{\eta}$ and $x^{*}$ may be far in a direction orthogonal to $c$ depending on $x_{0}$.
Example 2.6.
Figure 2 shows an example where $S\subseteq\mathbb{R}^{2}$ is a square centered at the origin with vertices $\{(\pm 1,\pm 1)\}$. Let $x_{0}=O$. For any $\eta>1$, there exists $c$ such that
$\|x^{*}-x^{\eta}\|^{2}\geq\frac{1}{4}.$
Indeed, let $c=(\tfrac{1}{2\eta},1)$. The optimal solution is $x^{*}=(1,1)$ while $x^{\eta}=(\tfrac{1}{2},1)$.
Figure 2. Illustration of Example 2.6. By selecting $\eta$ sufficiently large we can ensure $\langle c,x^{*}\rangle-\langle c,x^{\eta}\rangle\leq\epsilon$ independent of $c$, while $\|x^{*}-x^{\eta}\|$ remains large.
Now we show $x^{\eta}\to x^{*}$ when $\eta\to\infty$.
Note that both $x^{\eta}$ and $x^{*}$ are orthogonal projections, with $x^{\eta}=P_{S}(x_{0}+\eta c)$ and $x^{*}=P_{\mathcal{M}(c)}(x_{0})$.
Theorem 2.7.
Let $S\subseteq\mathcal{H}$ be a closed and convex set and $c\in\mathcal{H}$ such that $\mathcal{M}(c)$ is nonempty. Then
$\lim_{\eta\to\infty}P_{S}(x_{0}+\eta c)=P_{\mathcal{M}(c)}(x_{0}).$
Proof.
Lemma 2.4 implies that
$\|x^{\eta}-x_{0}\|\leq\|x^{*}-x_{0}\|$
for all $\eta>0$. Therefore, $x^{\eta}$ lies in a closed ball of radius $\|x^{*}\|$ centered at $x_{0}$. Closed balls in Hilbert space are weakly sequentially compact. Consider an arbitrary increasing sequence $\{\eta_{j}\}_{j\in\mathbb{N}}$ such that $x^{\eta_{j}}\rightharpoonup\overline{x}$ for some $\overline{x}\in S$. We show that $x^{\eta_{j}}\to x^{*}$ as $j\to\infty$.
Lemma 2.4 implies that $\overline{x}\in\mathcal{M}(c)$. Lemma 2.4 also implies that $\|\overline{x}-x_{0}\|\leq\|x^{*}-x_{0}\|$. Since $x^{*}$ is the unique element of $\mathcal{M}(c)$ closest to $x_{0}$, we conclude $\overline{x}=x^{*}$.
Since $\mathcal{H}$ is a real Hilbert space, it satisfies the Radon-Riesz property (see, e.g. [16]). To show strong convergence of $\{x^{\eta_{j}}\}$, it then remains to show that $\|x^{\eta_{j}}\|\to\|x^{*}\|$.
Lemma 2.4 implies $\|x^{\eta}-x_{0}\|^{2}\leq\|x^{*}-x_{0}\|^{2}$. Therefore $\limsup_{j}\|x^{\eta_{j}}-x_{0}\|^{2}\leq\|x^{*}-x_{0}\|^{2}$. The weak lower semicontinuity of the norm yields $\|x^{*}-x_{0}\|\leq\liminf_{j}\|x^{\eta_{j}}-x_{0}\|$. Since the limit inferior is at most the limit superior, combining both inequalities, we obtain
$\|x^{*}-x_{0}\|^{2}\leq\liminf_{j}\|x^{\eta_{j}}-x_{0}\|^{2}\leq\limsup_{j}\|x^{\eta_{j}}-x_{0}\|^{2}\leq\|x^{*}-x_{0}\|^{2}.$
Therefore, $\lim_{j}\|x^{\eta_{j}}-x_{0}\|^{2}=\|x^{*}-x_{0}\|^{2}$. Expanding both sides, we conclude that $\|x^{\eta_{j}}\|\to\|x^{*}\|$.
Since every accumulation point strongly converges to $x^{*}$, the full limit holds. ∎
One naturally expects the linear objective to grow with the scale factor $\eta$, so that larger $\eta$ yields progressively better solutions. Indeed, this monotonicity holds.
Proposition 2.8.
If $\eta_{1}<\eta_{2}$, then
$\langle c,x^{\eta_{1}}\rangle\leq\langle c,x^{\eta_{2}}\rangle.$
Proof.
By definition of $x^{\eta}$,
$\|(x_{0}+\eta_{1}c)-x^{\eta_{1}}\|^{2}\leq\|(x_{0}+\eta_{1}c)-x^{\eta_{2}}\|^{2}\text{ and }\|(x_{0}+\eta_{2}c)-x^{\eta_{2}}\|^{2}\leq\|(x_{0}+\eta_{2}c)-x^{\eta_{1}}\|^{2}.$
Expanding then adding the two inequalities, we obtain
$({\eta_{2}}-\eta_{1})\langle c,x^{\eta_{1}}\rangle\leq({\eta_{2}}-\eta_{1})\langle c,x^{\eta_{2}}\rangle.$
Since ${\eta_{2}}-\eta_{1}>0$, we obtain the desired inequality. ∎
3. Projected Subgradient Ascent
In this section, we analyze the behavior of projected subgradient ascent for maximizing a continuous convex function $f$ over a closed convex set $S\subseteq\mathcal{H}$. The problem of maximizing a convex function is NP‐hard (e.g., [20]). Therefore, we focus on obtaining first-order stationary points.
For convex minimization, projected subgradient descent (PGD) with a diminishing step size sequence $\{\eta_{k}\}_{k\in\mathbb{N}}$ satisfying
[…]
converges to a global minimizer (see, e.g., [7]). In contrast, when maximizing a convex function using projected subgradient ascent (PGA) convergence to a global optimum is not guaranteed, regardless of the choice of step sizes. Rather than vanishing steps, we focus on the case of arbitrarily large step sizes and show that PGA always converges to a first-order stationary point of $f$.
Definition 3.1.
(Projected Subgradient Descent) Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be a convex function, $S\subseteq\mathcal{H}$ be a nonempty closed convex subset, and $\{\eta_{k}\}_{k\in\mathbb{N}}$ be a sequence of step sizes.
Projected subgradient descent generates a sequence of iterates $\{x_{k}\}_{k\in\mathbb{N}}$ in $S$,
$x_{k+1}=P_{S}(x_{k}-\eta_{k}g_{k})$
where $g_{k}\in\partial f(x_{k})$ is any subgradient of $f$ at $x_{k}$.
Definition 3.2.
(Projected Subgradient Ascent) Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be a convex function, $S\subseteq\mathcal{H}$ be a nonempty closed convex subset, and $\{\eta_{k}\}_{k\in\mathbb{N}}$ be a sequence of step sizes.
Projected subgradient ascescent generates a sequence of iterates $\{x_{k}\}_{k\in\mathbb{N}}$ in $S$,
$x_{k+1}=P_{S}(x_{k}+\eta_{k}g_{k})$
where $g_{k}\in\partial f(x_{k})$ is any subgradient of $f$ at $x_{k}$.
Proposition 3.9 shows that in the convex maximization regime with PGA, the sequence $\{f(x_{k})\}$ is always nondecreasing. When $f$ is bounded, this implies $\{f(x_{k})\}$ converges. If the step sizes are bounded from above, Theorem 3.12 shows that the accumulation set of $\{x_{k}\}$ is connected. Moreover, if the step sizes are bounded from below, Theorem 3.14 shows all of the accumulation points of $\{x_{k}\}$ are first-order stationary points of $f$.
These results illustrate a contrast between convex minimization with PGD and convex maximization with PGA. The convex minimization setting requires vanishing steps for convergence, while in the convex maximization setting, convergence is guaranteed with large steps. In Section 3.3, we consider the limiting behavior when all of the step sizes go to infinity.
3.1. Linear Functions
Here, we consider the case of maximizing linear functions as it already captures the intuition that one can take large step sizes. The case of maximizing convex functions is a natural generalization.
For a linear objective $f(x)=\langle c,x\rangle$, maximizing $f$ with PGA is equivalent to minimizing f-f with PGD. In the finite-dimensional case, classical results using $L$-Lipschitz gradients guarantee both PGD and PGA converge with any constant non-zero step size (see, e.g., [6]).
In the general case of possibly infinite-dimensional spaces, weak convergence under arbitrary step sizes can be shown using notions from monotone operator theory (see, e.g., [5]). We formally state this known result (see, e.g., Theorem 2.5 in [4]) as Theorem 3.7 to build the intuition for taking large step sizes before we tackle the broader class of general convex functions.
Definition 3.3.
(Operator) An operator $T$ on a real Hilbert space $\mathcal{H}$ is a mapping
$T\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\rightrightarrows\mathcal{H}$,
meaning that each $x\in\mathcal{H}$ is assigned a (possibly empty) subset $T(x)\subseteq\mathcal{H}$. Its domain is
$\mathrm{dom}\,T=\{x\in\mathcal{H}\mathrel{\mathop{\ordinarycolon}}T(x)\neq\emptyset\},$
its range is
$\mathrm{ran}\,T=\{u\in\mathcal{H}\mathrel{\mathop{\ordinarycolon}}\exists x\in\mathcal{H}\text{ with }u\in T(x)\},$
and its graph is
$\mathrm{gra}\,T=\{(x,u)\in\mathcal{H}\times\mathcal{H}\mathrel{\mathop{\ordinarycolon}}u\in T(x)\}.$
Definition 3.4.
(Monotone Operator) An operator $T\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\rightrightarrows\mathcal{H}$ is called monotone if
$\langle x-y,u-v\rangle\geq 0\text{ for all }(x,u),(y,v)\in\mathrm{gra}\,T.$
Definition 3.5.
(Maximal Monotone Operator) A monotone operator $T\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\rightrightarrows\mathcal{H}$ is maximal monotone if its graph cannot be strictly enlarged without losing monotonicity.
Definition 3.6.
(Resolvent) Let $T\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\rightrightarrows\mathcal{H}$ be a maximal monotone operator and let $\lambda>0$. The resolvent of $T$ with parameter $\lambda$ is the (single-valued) mapping
$J_{\lambda T}=(\mathrm{Id}+\lambda T)^{-1}.$
Theorem 3.7.
Let $\{\eta_{k}\}_{k\in\mathbb{N}}$ be a sequence of step sizes such that $\sum_{k=0}^{\infty}\eta_{k}=\infty$. Let $\{x_{k}\}_{k\in\mathbb{N}}$ be a sequence of iterates generated by PGA with $f(x)=\langle c,x\rangle$. The sequence $\{x_{k}\}_{k\in\mathbb{N}}$ converges weakly to a point in $\mathcal{M}(c)$.
Proof.
Consider the operator
$A=N_{S}-c,$
which is maximal monotone because both $N_{S}$ and the constant operator are maximal monotone, and the interior of the domain of the constant operator is the entire Hilbert space (see, e.g., [21]).
Let $J_{\eta_{k}A}$ be the resolvent of $\eta_{k}A$ where $\eta_{k}>0$. Then, we observe that
$x_{k+1}=P_{S}(x_{k}+\eta_{k}c)=J_{\eta_{k}A}(x_{k})$
through the following sequence of equivalences:
$y=J_{\eta_{k}A}(x)$ $\Leftrightarrow y=(\mathrm{Id}+\eta_{k}(N_{S}-c))^{-1}(x)$
$\Leftrightarrow y=(\mathrm{Id}+N_{S}-\eta_{k}c)^{-1}(x)$
$\Leftrightarrow x\in y+N_{S}(y)-\eta_{k}c$
$\Leftrightarrow x+\eta_{k}c-y\in N_{S}(y)$
(by Lemma 2.3) $\Leftrightarrow y=P_{S}(x+\eta_{k}c).$
By assumption, $\mathcal{M}(c)\neq\emptyset$. We observe that the zeros of $A$ equals $\mathcal{M}(c)$,
$\mathrm{zer}\,A=\{x\in\mathcal{H}\mathrel{\mathop{\ordinarycolon}}0\in Ax\}=\{x\in\mathcal{H}\mathrel{\mathop{\ordinarycolon}}c\in N_{S}(x)\}=\mathcal{M}(c).$
Since $A$ is a maximally monotone operator such that $\mathrm{zer}\,A\neq\emptyset$, and $x_{k+1}=J_{\eta_{k}A}(x_{k})$, we conclude that the sequence converges weakly to a point in $\mathrm{zer}\,A=\mathcal{M}(c)$ (see, e.g., Theorem 23.41 of [5]). ∎
Remark 3.8.
We emphasize the difference between the results in Theorem 2.7 and Theorem 3.7. The first theorem analyzes the behavior of a single step of projected gradient ascent $P_{S}(x_{0}+\eta c)$ as the step size $\eta$ goes to infinity, proving strong convergence to the unique solution in $\mathcal{M}(c)$ closest to $x_{0}$. In contrast, the second theorem establishes weak convergence of projected gradient ascent to some solution in $\mathcal{M}(c)$ after an infinite number of finite steps.
3.2. Convex Functions
Now, we consider the case of maximizing a continuous convex function with PGA.
In Theorem 3.14, we establish first-order stationarity of accumulation points under weak assumptions. We consider the case of general Hilbert spaces and only require that the step sizes do not vanish and remain bounded above. Neither differentiability nor Lipschitz continuity of the gradient is required.
Previous results that apply to the setting of maximizing convex functions require significantly stronger assumptions. For example, the convergence results in [2] apply in finite-dimensional spaces when $f$ is differentiable with a Lipschitz continuous gradient over a nonempty closed set $S$, and satisfies the Kurdyka-Łojasiewicz property.
Proposition 3.9.
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be convex, $S\subseteq\mathcal{H}$ be a convex set and $\{x_{k}\}_{\mathbb{N}}$ be a sequence of iterates generated by PGA. Then the sequence of values $\{f(x_{k})\}_{k\in\mathbb{N}}$ is nondecreasing.
Proof.
Let
$f_{k}(x)=f(x_{k})+\langle x-x_{k},g_{k}\rangle.$
Since $f$ is convex, and $g_{k}$ is a subgradient of $f$ at $x_{k}$, we know $f_{k}(x)$ lower-bounds $f(x)$, and they touch at $x_{k}$. In other words,
  1. (a)
    $f_{k}(x)\leq f(x)$ for all $x\in\mathcal{H}$, and
  2. (b)
    $f_{k}(x_{k})=f(x_{k})$.
We first show that
(2) $f_{k}(x_{k+1})\geq f_{k}(x_{k}).$
Note that $x_{k+1}=P_{S}(x_{k}+\eta_{k}g_{k})$ and $x_{k}=P_{S}(x_{k})$. By Proposition 2.8 with $\eta_{1}=0$ and $\eta_{2}=\eta_{k}$ we know that $\langle g_{k},x_{k+1}\rangle\geq\langle g_{k},x_{k}\rangle$. This implies $\langle x_{k+1}-x_{k},g_{k}\rangle\geq 0$ and now (2) follows from the definition of $f_{k}$.
The nondecreasing property follows from (a), (2), and (b),
$f(x_{k+1})\geq f_{k}(x_{k+1})\geq f_{k}(x_{k})=f(x_{k}).$
Corollary 3.10.
If $S$ is bounded, the sequence $\{f(x_{k})\}$ converges.
Proof.
By Proposition 3.9, the sequence is nondecreasing. Because $S$ is bounded, the sequence is also bounded. Finally, bounded nondecreasing sequences in $\mathbb{R}$ converge. ∎
Observation 3.11.
If $S$ is compact then $\{x_{k}\}$ has at least one accumulation point.
Theorem 3.12.
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be continuous and convex, $S\subseteq\mathcal{H}$ be a compact and convex set, and $\{x_{k}\}_{\mathbb{N}}$ be a sequence generated by PGA with step sizes $\{\eta_{k}\}_{\mathbb{N}}$ where $\limsup_{k\to\infty}\eta_{k}<\infty$. Then the set of accumulation points of $\{x_{k}\}_{\mathbb{N}}$ is connected.
Proof.
Let $y_{k}=x_{k}+\eta_{k}g_{k},v_{k}=x_{k+1}-x_{k}$, and $\Delta_{k}=f(x_{k+1})-f(x_{k})$. Then $x_{k+1}=P_{S}(y_{k})$.
We show that $\|v_{k}\|\to 0$ as $k\to\infty$.
By the optimality conditions of projections,
$\langle\eta_{k}g_{k}-v_{k},-v_{k}\rangle=\langle y_{k}-x_{k+1},x_{k}-x_{k+1}\rangle\leq 0,$
Therefore,
(3) $\eta_{k}\langle g_{k},v_{k}\rangle\geq\|v_{k}\|^{2}.$
Next, by the convexity of $f$,
(4) $f(x_{k+1})\geq f(x_{k})+\langle g_{k},v_{k}\rangle.$
Combining inequalities (3) and (4), we obtain
$\eta_{k}\Delta_{k}\geq\|v_{k}\|^{2}.$
Let $\overline{\eta}=\lim\sup_{k\to\infty}\eta_{k}$. There exists some $K\in\mathbb{N}$ such that for any $k\geq K$, $\overline{\eta}+1>\eta_{k}$. Hence, for $k\geq K$,
$(\overline{\eta}+1)\Delta_{k}>\eta_{k}\Delta_{k}\geq\|v_{k}\|^{2}.$
By Corollary 3.10, we know $(\overline{\eta}+1)\Delta_{k}\to 0$, implying that
$\|x_{k+1}-x_{k}\|^{2}=\|v_{k}\|^{2}\to 0.$
Therefore, the set of accumulation points is connected (see, e.g., [3]). ∎
Definition 3.13 (First-Order Stationarity).
A point $x\in S$ is a first-order stationary point for the maximization of $f$ over $S$ if there exists a subgradient $g\in\partial f(x)$ such that $\langle g,z-x\rangle\leq 0$ for all $z\in S$. Equivalently, if there exists $g\in\partial f(x)$ such that $g\in N_{S}(x)$.
Theorem 3.14.
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be continuous and convex, $S\subseteq\mathcal{H}$ be a compact and convex set, and $\{x_{k}\}_{\mathbb{N}}$ be a sequence generated by PGA with step sizes $\{\eta_{k}\}_{\mathbb{N}}$ where $\limsup_{k\to\infty}\eta_{k}<\infty$ and $\liminf_{k\to\infty}\eta_{k}>0$. Then every accumulation point of $\{x_{k}\}_{\mathbb{N}}$ is a first-order stationary point of $f$.
Proof.
[…]
Let $\underline{\eta}=\liminf_{k\to\infty}\eta_{k}>0$. Then there exists $K\in\mathbb{N}$ such that for any $k\geq K$, $\underline{\eta}/2<\eta_{k}$.
Let $v_{k}=x_{k+1}-x_{k}$. The proof of Theorem 3.12 showed $\|v_{k}\|\to 0$. Therefore,
$\lim_{k\to\infty}\frac{\|v_{k}\|}{\eta_{k}}\leq\lim_{k\to\infty}\frac{2\|v_{k}\|}{\underline{\eta}}=0.$
Since $\frac{v_{k}}{\eta_{k}}$ converges to 0 in norm,
(5) $\lim_{k\to\infty}\frac{v_{k}}{\eta_{k}}=0.$
Because $\tilde{x}$ is an accumulation point, there exists a subsequence $\{x_{k_{j}}\}$ such that $x_{k_{j}}\to\tilde{x}$. Additionally, $\|v_{k_{j}}\|\to 0$ implies that $x_{k_{j}+1}\to\tilde{x}$.
Because $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ is convex and continuous, for any $x\in\mathcal{H}$, there exists $r_{x}>0$ and $L_{x}$ such that $\partial f(B(x,r_{x}))\subset B(0,L_{x})$ (see, e.g., Proposition 16.17 of [5]). The balls form an open cover of $S$. Compactness of $S$ yields a finite subcover, with which we may find some $L$ such that $\{g_{k_{j}}\}\subseteq\bigcup_{x\in S}\partial f(x)\subseteq\overline{B}(0,L).$ Because $\overline{B}(0,L)$ is weakly compact, there exists a further subsequence $\{g_{k_{j_{\ell}}}\}$ such that $g_{k_{j_{\ell}}}\rightharpoonup g$ for some $g$. Combining with (5), $g_{k_{j_{\ell}}}-{v_{k_{j_{\ell}}}}/{\eta_{k_{j_{\ell}}}}\rightharpoonup g$.
[…]
3.3. Conditional Gradient and Iterated Linear Optimization
Now we consider the limiting case of PGA when all of the step sizes go to infinity and relate this limit to the conditional gradient method and iterated linear optimization.
Consider the limit of the $k$-th PGA iteration as $\eta_{k}\to\infty$,
$x_{k+1}=\lim_{\eta_{k}\to\infty}P_{S}(x_{k}+\eta_{k}g_{k}).$
When $f$ is differentiable $g_{k}=\nabla f(x_{k})$ and by Theorem 2.7,
(6) $x_{k+1}=P_{\mathcal{M}(\nabla f(x_{k}))}(x_{k}).$
That is, $x_{k+1}$ is the maximizer of $\langle\nabla f(x_{k}),x\rangle$ closest to $x_{k}$.
This limiting behavior of PGA with infinity step size is closely related to the conditional gradient method, also known as the Frank-Wolfe algorithm [14]. This parallels the convex minimization setting, where the limit of a PGD step in a polytope is known to recover a solution of the corresponding linear minimization problem [17].
Definition 3.15.
(Conditional Gradient/Frank-Wolfe) Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{H}\to\mathbb{R}$ be convex and differentiable, $S\subseteq\mathcal{H}$ be nonempty, closed, and convex, and $\{\eta_{k}\}_{k\in\mathbb{N}}\subseteq[0,1]$ be a sequence of step sizes, The CG algorithm generates a sequence of iterates,
$x_{k+1}=x_{k}+\eta_{k}(z_{k}-x_{k}),$
where
$z_{k}\in\operatorname*{\arg\!\max}_{z\in S}\,\langle\nabla f(x_{k}),z\rangle.$
The unit-step variant of conditional gradient (CGU) sets $\eta_{k}=1$. This yields
$x_{k+1}\in\operatorname*{\arg\!\max}_{z\in S}\,\langle\nabla f(x_{k}),z\rangle=\mathcal{M}(\nabla f(x_{k})).$
Now we can see that PGA with infinite step sizes, as defined by Equation (6), is a deterministic variant of CGU, where in each iteration we select the particular element of $\mathcal{M}(\nabla f(x_{k}))$ that is closest in norm to the last iterate. When $|\mathcal{M}(\nabla f(x_{k}))|=1$, such as when $S$ is smooth, the methods coincide.
Finally, we note that when $f(x)=\tfrac{1}{2}\|x\|^{2}$ the CGU iteration leads to,
$x_{k+1}\in\operatorname*{\arg\!\max}_{z\in S}\,\langle x_{k},z\rangle,$
which is exactly the update rule defined by the iterated linear optimization paradigm described in [13]. That is, iterated linear optimization is equivalent to CGU with a particular choice for $f$, and PGA with infinite step size defines a deterministic variant of both methods.
Acknowledgements
This research was partially funded by the Brown University Advanced Undergraduate Research Fellowship under the SPRINT/UTRA program.